OpenStudy (anonymous):

True or False? The volume of a cone with a base AREA of 12 cm squared and a height of 9 cm is 72 to the third power.

jimthompson5910 (jim_thompson5910):

hint: Volume of a Cone V = (1/3)*B*h B = area of base h = height

OpenStudy (agent0smith):

Volume of a cone or pyramid:\[\Large V = \frac{ 1 }{ 3 }B *h\]

OpenStudy (agent0smith):

B = 12 cm^2 and h=9cm

OpenStudy (anonymous):

Thanks guys, the answer is false, then, right?

jimthompson5910 (jim_thompson5910):

yes that's correct, you should get 36 cm^3 and not 72 cm^3